### v1

# MuJo: Multimodal Joint Feature Space Learning for Human Activity Recognition

Abstract
Human Activity Recognition is a longstanding problem in AI with applications in a broad range of areas: from healthcare, sports and fitness, security, and human computer interaction to robotics. The performance of HAR in real-world settings is strongly dependent on the type and quality of the input signal that can be acquired. Given an unobstructed, high-quality camera view of a scene, computer vision systems, in particular in conjunction with foundational models (e.g., CLIP), can today fairly reliably distinguish complex activities. On the other hand, recognition using modalities such as wearable sensors (which are often more broadly available, e.g, in mobile phones and smartwatches) is a more difficult problem, as the signals often contain less information and labeled training data is more difficult to acquire. In this work, we show how we can improve HAR performance across different modalities using multimodal contrastive pretraining. Our approach MuJo (Multimodal Joint Feature Space Learning), learns a multimodal joint feature space with video, language, pose, and IMU sensor data. The proposed approach combines contrastive and multitask learning methods and analyzes different multitasking strategies for learning a compact shared representation. A large dataset with parallel video, language, pose, and sensor data points is also introduced to support the research, along with an analysis of the robustness of the multimodal joint space for modal-incomplete and low-resource data. On the MM-Fit dataset, our model achieves an impressive Macro -Score of up to 0.992 with only 2% of the train data and 0.999 when using all available training data for classification tasks. Moreover, in the scenario where the MM-Fit dataset is unseen, we demonstrate a generalization performance of up to 0.638.
1 Introduction
Perceiving and interpreting human activity is a core functionality of intelligent systems with applications in a broad range of areas: from healthcare Tan et al. (2021) through sports and fitness Host and Ivašić-Kos (2022); Nadeem et al. (2020), security Sunil et al. (2021) to robotics Piyathilaka and Kodagoda (2015). A basic capability needed to facilitate such functionality is the identification of concrete human actions from the system’s sensory input, referred to as Human Activity Recognition (HAR) Hamad et al. (2023). In general, HAR is a difficult problem, especially when applied to uncontrolled real-life environments, where it must contend with the variability of human actions and the open-ended variations in environmental conditions. In such settings, the performance of HAR systems is strongly dependent on the available input modalities. Recent advances in Computer Vision, particularly with Large Vision Language Models (VLMs) Radford et al. (2021); Achiam et al. (2023), have enabled impressive performance when given sufficiently high-quality images. In many cases, merely feeding images to publicly available tools such as GPT-4 can result in accurate description of scenes involving complex and subtle activities. However, in real-life settings, such images are not always available; either due to a lack of cameras or because of privacy concerns. Much more broadly available input modalities are wearable sensors, in particular, Inertial Measurement Units (IMUs), which are ubiquitous in smartphones, smartwatches or even earbuds, all of which people carry with them on a daily basis. Unfortunately, achieving acceptable recognition performance with such sensors is a much more difficult problem due to two factors:
1.
Information ambiguity: An IMU (which is essentially a motion sensor) in a smartwatch on the wrist or in a smartphone in a pocket contains much less detailed information about the user’s activity than a high-quality picture.
2.
Lack of training data: The success of VLMs is built on the online availability of massive amounts of text/image data pairs, which can be used for training aligned representations. Although large amounts of unlabeled IMU data are increasingly becoming available Chan et al. (2024), labeled training data remains scarce.
The work presented in this paper, which aims to address the issues outlined previously, is founded on three key observations:
1.
While the input modalities available during inference—the time when the system is utilized in the real world—may be constrained, the input modalities during training can be selected with much greater flexibility.
2.
Representation learning facilitates knowledge transfer between modalities by aligning embedding spaces, in particular when synchronized training data is available.
3.
Significant progress has recently been made in generating synthetic data for a variety of sensors from videos. For example, IMUTube Kwon et al. (2021) generates virtual gyroscope and accelerometer data from videos, while pose extraction methods like OpenPose Cao et al. (2017) and VideoPose3D Pavllo et al. (2019) are used to generate skeletal data of human activities. This enables the creation of synchronised video/sensor training data and through the alignment of video with text, it facilitates the synchronization of sensor data with textual descriptions.
Based on the above considerations, we propose a multimodal contrastive pretraining method that creates a joint representation between videos, video-derived poses, synthetic IMU data created from videos using IMUTube Kwon et al. (2021), and textual descriptions. Our approach, which we refer to as MuJo, aims to utilize the complementary information of each modality to improve the performance for both unimodal and multimodal inference. In particular, we show that our method dramatically reduces the need for labeled training data for downstream tasks. Thus, on the well-known MM-Fit dataset Strömbäck et al. (2020), when using all available training data, our method improves the recognition rate (Macro -Score) from 0.973 to 0.999 in the multimodal case (encompassing all of the above modalities), from 0.957 to 0.988 for poses, and from 0.727 to 0.863 for video alone. For sensor data, the performance of our method is slightly worse (0.798 as opposed to 0.822), which is due to the fact that MM-Fit uses real, rather than simulated sensor data. However, when using just 2% of the available data for training, the baseline with sensors has an -Score of just 0.378 whereas our systems achieves a reasonable 0.627 (see section 6.2 for details). The multimodal -Score is 0.819 in the baseline, while our method raises it to 0.992.
Overall, the key contributions of this work can be summarized as follows:
Development of a large fitness dataset111Available after publication at link. with parallel video, language (i.e., labels of video clips, captions uttered by the instructor in the video clip, and precise descriptions of the exercises generated by GPT-3.5), pose, and simulated IMU data points.
Learning of an effective joint space for the modalities of text, video, pose, and sensor by leveraging the contrastive loss.
Analysis of usefulness and generalizability of multimodal joint space for each modality and their combination, as well as the benefits of pretraining when a low amount of data is available for the target task.
2 Related Work
2.1 Human Activity Recognition
Unimodal activity recognition.
Unimodal activity recognition has been extensively studied in HAR. Various approaches have been proposed to address this task, and they can be broadly categorized into two groups: hand-crafted feature-based and deep learning-based methods. Hand-crafted feature-based methods rely on extracting low-level features from the input data and designing a classifier to recognize activities. For instance, in accelerometer-based activity recognition Kwapisz et al. (2011), features such as mean, standard deviation, and energy of the acceleration signal are often used to capture the characteristics of different activities. Popular classifiers used in this approach include decision trees, support vector machines, and k-nearest neighbors. Deep learning-based methods, on the other hand, use neural networks to learn features and recognize activities from raw input data. Convolutional neural networks (CNNs) and recurrent neural networks (RNNs) are two popular types of neural networks used in this approach. CNNs are commonly used for image-based activity recognition Strömbäck et al. (2020); Ordóñez and Roggen (2016), being combined with RNNs in some cases Ordóñez and Roggen (2016). Despite the success of unimodal activity recognition, it is often limited by the fact that a single modality of input data may only capture some of the necessary information about the activity, leading to lower recognition accuracy. Therefore, multimodal activity recognition, which combines information from multiple sources, has gained increasing attention in recent years.
Multimodal activity recognition.
Multimodal activity recognition recognizes human activities using information from multiple modalities, such as IMU data, audio, and video Mekruksavanich and Jitpattanakul (2022); Ijaz et al. (2022). These features can then be combined using fusion methods such as early fusion or late fusion Duhme et al. (2022). Some studies have also focused on domain adaptation, where the goal is to recognize activities in a new environment using data from a different environment Plananamente et al. (2022). In such cases, transfer learning techniques can adapt the model to the new environment using a small amount of labeled data. Overall, multimodal activity recognition has shown promising results, with improved accuracy compared to unimodal approaches. However, challenges in data fusion, feature extraction, and model selection still need to be addressed to improve the performance of multimodal activity recognition systems Islam et al. (2022).
2.2 Representation Learning for Human Activities
Representation learning with proxy tasks.
Different modalities, e.g., natural language, visual inputs, or sensor signals, often complement a common concept. A key challenge of multimodal feature representation is exploring efficient methods for multimodal fusion with the given modalities to avoid missing modal features. We can differentiate the representation learning methods in two aspects: learning objectives and fusing the learned features in a joint space. In learning objectives, various approaches have been proposed to learn multimodal representations by formulating proxy tasks such as reconstruction, completion Seo et al. (2021); Sun et al. (2019), matching/alignment Li et al. (2020); Miech et al. (2020), and ordering tasks Li et al. (2020). Generating one feature vector with multimodal inputs for a common concept is a crucial problem of multimodal representation learning. Simple methods like concatenating Hu et al. (2018); Strömbäck et al. (2020) learned uni-modal feature vectors have been extensively explored to enhance the performance of downstream tasks. These methods involve combining feature vectors from a single modality, such as text or image, to create more informative representations that can be used for various tasks. Thereafter, Zadeh et al. proposed a tensor fusion network in Zadeh et al. (2017). Lu et al. proposed a co-attention mechanism based on the transformer architecture in Lu et al. (2019) for efficient cross-modality learning.
Representation via contrastive learning.
Contrastive Learning is as an efficient self-supervised framework applied across multiple domains, which learns similar/dissimilar representations from data that are organized into similar/dissimilar pairs. Recently, Radford et al. presented a novel approach, known as CLIP Radford et al. (2021), to jointly represent images and text descriptions by leveraging contrastive loss with the similarity of image-text pairs. CLIP learns two feature spaces with two encoders, i.e., image and text encoders, then projects the learned features into a shared latent space with contrastive loss. Contrastive learning, a joint feature space of text-image pairs, has been extended to other modalities. For example, video-text data has attracted studies on shared space of video-text features Xu et al. (2021); Xue et al. (2022), text-to-video generation Singer et al. (2022). Thereafter, MotionCLIP Tevet et al. (2022) trained an encoder to find the proper embedding of an input sequence in CLIP space, and a decoder that generates the most fitting motion to a given CLIP space. Following that, IMU2CLIP Moon et al. (2022) has been proposed to learn a joint space to align Inertial Measurement Unit (IMU) motion sensor recordings with video and text, by projecting them into the joint representation space.
Studies in action recognition Aggarwal and Ryoo (2011); Li et al. (2020); Wang et al. (2021); Sun et al. (2022); Morshed et al. (2023) can also significantly enhance the field of recognition in fitness activities, despite action recognition emphasis on a broad spectrum of actions such as “cutting the potato” and “cleaning the drawer”, among others. Certain investigations in action recognition also endeavor to acquire knowledge of the feature space through contrastive learning methodologies, exemplified by initiatives such as HERO Li et al. (2020) and ActionCLIP Wang et al. (2021).
The achievements discussed in the aforementioned studies have shown contrastive learning has a high impact on extracting a joint space with similar and dissimilar pairs. Hence, the contrastive learning methodology employed in this study is utilized to acquire knowledge of the feature space pertaining to fitness activities. Contrary to previous work, we extend this line of work on contrastive learning to a unique multimodal setting that utilizes more than two modalities, such as video, IMU sensors, poses, and text, within a multitask training manner in an end-to-end training paradigm similar to ImageBind Girdhar et al. (2023). Thus, our model and data provide a cross-modal classification opportunity.
3 Dataset
3.1 Video Collection
YouTube provides chapters within a video, each with its preview, providing information and context for different parts of the video. Video owner can manually add chapters to their uploaded videos or rely on automatic chapter generation. We manually collect 1212 fitness videos with pre-annotated video chapters. Additionally, we retrieve the captions automatically generated by YouTube. Our videos include only one tutor who teaches how to do fitness exercises (e.g., “squat with dumbbells”, “sit-ups”, “side deep squats”, etc.). Thus, the exercise videos provide instructions for the activities while showing the activities clearly. The annotation does not include predefined categories or labels. Each video owner annotates the activity segment of the videos with their personal preferences, such as “deep squats with dumbbells”, “high jump squats”, etc., instead of only “squats”. Given the temporal boundaries of the activity within each video, we split the videos in shorter activity clips along with their corresponding chapters and caption segments.
Since workout exercises are typically short, activity videos that are too long often contain multiple exercises and, thus, incorrect chapter labels. Hence, we only keep exercises shorter than 2 minutes and discard more extended exercises. After that, the dataset consists of 10,695 (video, chapter, caption) samples.
3.2 Dataset Construction
We convert all videos to 720p resolution and 50 frames per second via FFmpegs adaptive overlapped block motion compensation with bidirectional motion estimation. Then, we extract simulated sensor data, poses, as well as a variety of video and text features. Video features are extracted with HERO Li et al. (2020) utilizing CLIP Radford et al. (2021) and S3D Xie et al. (2018) embeddings. On the textual side, we apply several normalization strategies to the raw chapters such as replacing abbreviations within the text (e.g., changing “l” to “left” and “r” to “right”) as well as removing stop words, numbers, and a few special symbols from a predefined list. Further, we use GPT 3.5 to generate detailed descriptions of the exercises. We then extract multiple text features with different models, including OpenAI’s text-embedding-ada-002 embedding model (ADA), CLIP, BERT Devlin et al. (2018), as well as language features and motion tokens extracted with Text2Motion Transformer Guo et al. (2022). We apply IMUTube Kwon et al. (2021) to generate virtual gyroscope and accelerometer data for sensors placed on all 16 joint nodes. We also use the implementation of IMUTube to extract 3D poses from the videos, which utilizes OpenPose followed by VideoPose3D for 3d pose lifting.
Since IMUTube fails to extract sensor data for 884 videos, we exclude these sequences. We use the remaining 9811 exercises as our dataset and split it into train (8,518 videos), validation (946 videos) and test (347 videos) sets. Since workout session videos typically start with an introductory sequence and conclude with a farewell, we discard the initial and final 15%222This value has proven to be effective in experimental investigations. of each video. Despite these filtering steps, we receive a large dataset totaling around 78 hours in length. Instances are generated using sliding windows of 100 frames in length (i.e., 2 seconds) with 50 frames overlap from the videos. We found this size suitable for fitness exercises as a 2 second window generally captures enough meaningful information regarding the subject’s movements.
In summary, our dataset comprises a large collection of short fitness activity clips containing video features, poses, synthetic IMU data, chapter labels, captions, and GPT-generated descriptions as well as several types of features extracted from these text modalities.
4 Method
Our method leverages multimodal information for HAR pretraining. The fundamental idea is that the extracted information for a short video interval should be similar, as it represents the same activity from various perspectives. For each window, we extract features for each modality, which means our data consists of where is the feature for modality at time interval . We use as modalities
: The sequence of 3D poses for the subject extracted using VideoPose3D.
: Simulated accelerometer data of both wrists generated by ImuTube Kwon et al. (2021).
: Embeddings extracted using OpenAI’s ADA model from YouTube’s chapter information with a lightweight normalization process removing a few special symbols and replacing abbreviations within the text.
Each modality has its own encoder network and projection layers . Thus, serves as our representation for a vector of modality in our shared representation space. We learn this shared space between modality representations using pair-wise contrastive learning, as depicted in Figure 1. First, for all possible modality pairs, we compute similarities using
Then, those similarities are softmax-normalized within the batch to calculate the loss between the two modalities
where and correspond to aligned batch data for modalities and , respectively, and is the temperature hyper-parameter. Since we calculate losses between all modalities, our total loss amounts to
where is the total number of modalities (in our case 4). An example can be seen in Figure 1.
Regarding the structure of our encoders, we utilize a model with a single fully connected layer for and two fully connected layers for . For and , we employ the CNN encoder architecture as described in Strömbäck et al. (2020). The projection is a two-layer fully connected model for all modalities , mapping them to feature vectors of length 1280.
We apply a random search approach for hyperparameter selection. Our hyperparameters include learning rate, number and size of projection layers, dropout rates, and activation functions. We also experimented with various configurations of the loss function, specifically comparing using the text modality as an anchor and, hence, considering sensor-text, pose-text, and video-text loss pairs only versus incorporating all possible modality combinations into the loss. Encompassing all modality pairs together yielded superior results. We explored weighting the text modality higher than the others in the loss, but it did not have any observable effect on performance.
We also explored the use of S3D-based video embeddings, the gyroscope as sensor input, and the other text modalities and features described in Section 3.2 as inputs. However, since those configurations did not yield the best performance during the hyper-parameter search, we have omitted those results for the sake of brevity.
|c|c|ccc|ccc|
Dataset Type Input Modality Mean Macro -Score
Train Data: Train Data:
Baseline Ours (Frozen) Ours Baseline Ours (Frozen) Ours
Real-World Sensor
Pose
Video
Multimodal
Rendered Sensor
Pose
Video
Multimodal
|c|c|c|ccc|ccc|
Input Modality Calibration Mean Macro -Score
Train Data: Train Data:
Baseline Ours (Frozen) Ours Baseline Ours (Frozen) Ours
Without “No Activity”
Sensor No
Unsupervised
Supervised
Pose No
Unsupervised
Supervised
Video No
Multimodal No
Unsupervised
Supervised
With “No Activity”
Sensor No
Unsupervised
Supervised
Pose No
Unsupervised
Supervised
Video No
Multimodal No
Unsupervised
Supervised
5 Experiments
We evaluate the performance of our methodology on the FLAG3D and MM-Fit datasets. On FLAG3D, we show performance in classification task whereas on MM-Fit, besides classification, we additionally evaluate the generalization performance of our representation in the scenario where the dataset is unseen and also compare unimodal versus multimodal training.
5.1 Classification
For classification, we utilize a classifier architecture comprising one or multiple pretrained encoder(s) , projection(s) , and a classification head consisting of 2 layers. We contrast three scenarios: In the baseline setting, all weights are initialized randomly, meaning the model lacks any pretrained knowledge. In the subsequent scenarios, the encoder and projection head is initialized with weights from our pretrained model. In one case, the weights of the encoder and projection head are fixed (frozen), whereas in the other case, we fine-tune them during the classifier’s training process. We train the model with all possible input modalities, namely accelerometer, poses, and videos, separately. Furthermore, we experiment with multimodal input by providing all three modalities to the model and concatenating the outputs from the respective projections before feeding them into . In addition, we train with different proportions of the training data, ranging from one to 100 percent, with the training data being sampled per class to ensure balanced representation. To guarantee comparability of all results, we train each classifier with the same settings: We utilize a learning rate of 0.0001 and training is conducted for a maximum of 200 epochs, incorporating early stopping after 25 epochs without improvement. We employ a weighted cross-entropy classification loss to address class imbalance, given by:
where is the number of classes, is a one-hot encoded ground truth activity label and is the predicted label vector of for input , that is, the application of our classification layers to obtain . The weight for class is calculated by dividing the number of data samples of the largest class by the number of instances in class itself. To ensure validity of the results, we train 20 models independently for each configuration, with the training data being randomly selected anew for each run (for proportions less than 100%) and the non-pretrained weights being reinitialized randomly.
5.1.1 FLAG3D
The FLAG3D Tang et al. (2023) dataset is a large-scale 3D fitness activity collection with language instruction of 60 categories. The dataset comprises (a) 7,200 3D activity sequences performed by 10 individuals acquired from an advanced MoCap system, (b) 172,800 2D RGB videos rendered from the 3D MoCap data including different scenes, avatars, and camera positions, and (c) 7,200 authentic videos of ten persons captured by cost-effective smartphones in both indoor and outdoor natural settings. Additionally, FLAG3D offers comprehensive and meticulously crafted sentence-level language instructions for each fitness activity.
We evaluate our method on all 7,200 real-world videos from Flag3D. To prepare the data, we follow the same approach as described in Section 3.2: First, we convert the videos to 50 fps. Then, we extract CLIP video features, ADA text embeddings, virtual sensor data, and poses (VideoPose3D) with IMUTube. We discard 943 videos for which IMUTube was unable to generate any sensor data. For the remaining 6,257 videos, we retrieve cumulatively over 21 hours of useful data. We split the data based on the person IDs performing the exercises into train (5 persons), validation (2 persons), and test (3 persons) sets. We apply a sliding window with a window size of 100 frames (2 seconds) and a step size of 50 frames (1 second). We also evaluate our method on a subset of 7,200 rendered videos, from which IMUTube was capable to extract sensor data for 6,832 videos (totaling 24 hours), processing them analogously to the real-world videos.
5.1.2 MM-Fit
The MM-Fit Strömbäck et al. (2020) dataset includes multi-view RGB-D video recordings, accompanied by 3D pose estimations extracted from single-view RGB frames with OpenPose and VideoPose3D. Additionally, it comprises a comprehensive compilation of time-synchronized inertial sensor data sourced from two smartphones, two smartwatches, and a earbud worn by participants during workout sessions. 10 Participants performed 21 full-body workout routines in total while positioned in front of two depth cameras, with five smart devices strategically placed to collect inertial sensor data. The task is delimited to the recognition of exercises performed by individual participants. Thus, each workout session involves a solitary participant. This setup is tailored for home workout environments. Each workout session encompasses three sets of 10 exercises, with 10 repetitions for each exercise. The selected exercises encompass well-known resistance training routines, including squats, lunges (with dumbbells), alternating-arm bicep curls, sit-ups, push-ups, seated overhead dumbbell triceps extensions, standing dumbbell rows, jumping jacks, seated dumbbell shoulder press, and dumbbell lateral shoulder raises.
We conduct the same preprocessing steps as for our dataset and FLAG3D but solely extract ADA text embeddings and CLIP video features. For pose and sensor, we utilize the 3D pose and Smartwatch accelerometer data of the left and right wrist provided by MM-Fit. Furthermore, we consider two scenarios: training the classifiers with and without including the “No Activity” category. Additionally, we calibrate sensor and pose data using a probability integral transform function Conover and Iman (1981), following the implementation of Kwon et al. (2021). In the unsupervised setting, all real data is calibrated with the virtual train data; whereas in the supervised case, label information is utilized to perform the calibration separately for each class.
5.2 Evaluation of Generalization Performance on Unseen Datasets
To evaluate performance in a more realistic scenario, we measure the model’s capability to match input modalities with the correct labels of the MM-Fit dataset in the scenario where the dataset is unseen. For that, we first calculate the ADA text embeddings for all labels of the MM-Fit dataset. For the “No Activity” class, which doesn’t correspond to a specific activity, we generate its embedding by taking the mean of the embeddings of all labels. For a given data sample of input modality , we then choose labels based on the closest similarity score by calculating for all labels .
6 Results
6.1 FLAG3D
Table 4 demonstrates that the pretrained model consistently outperforms the baseline across all input modalities with respect to Macro -Score on both the real-world and the rendered dataset when utilizing either 2% or 100% of the training data. When comparing the baseline to our unfrozen pretrained model across both datasets and all types of input, we observe a absolute performance increase ranging from 0.042 (rendered dataset, video input) to 0.219 (real-world, multimodal) with just 2% of the training data. For the full dataset (100% of the training data), the improvement spans from 0.029 (real-world, video) to 0.208 (real-world, multimodal). This indicates that our pretrained encoders have learned a robust representation of the respective input modality, thereby enhancing classification performance even on a new dataset. We expect that our pretrained, unfrozen encoders would outperform the frozen ones by adapting more effectively to the training data. This expectation was confirmed in 7 out of 8 scenarios when using the full dataset for training, with a absolute performance increase of 0.014 on average. However, with only 2% of the training data, the frozen encoders show even superior performance in 3 out of 8 scenarios, and the unfrozen encoders do not perform significantly better on average, suggesting that the pretrained representation could not be enhanced with such a limited amount of training data. Generally, poses perform best out of all input modalities. While the multimodal setting does not show substantial improvement for the rendered dataset, a slight enhancement (0.015 for 2% train and 0.028 for 100% train) is observed for our pretrained, unfrozen model in the real-world setting when concatenating all modalities, compared to just using poses as input.
From the boxplots presented in Figure 3, which compare the performance of the models across all modalities and various training fractions ranging from 1% to 100%, we observe that the pretrained models significantly outperform the baseline for all training fractions and most modalities. The only exception is the video modality, which performed worse out of all modalities and only slightly better than the baseline. However, with a limited amount of training data, our model demonstrates better performance relative to the baseline in the classification task over all modalities.
6.2 MM-Fit
Table 4 presents the results of experiments conducted with and without the inclusion of the “No Activity” class, across all input modalities and calibration configurations (none, unsupervised, and supervised), for both 2% and 100% training data fractions. We find that applying supervised, per class calibration generally enhances performance, whereas unsupervised calibration tends to diminish performance for sensor input. For pose data and multimodal inputs (which include poses), our pretrained encoders are unable to yield results without calibration—the training loss failed to decrease, remaining at infinity for 25 epochs. Notably, both IMUTube and MM-Fit utilize VidePose3D for pose extraction, but IMUTube incorporates extensive additional pre- and post-processing steps. Given that our pretraining utilized these poses generated by IMUTube, we speculate that our pose representation learned is significantly different from that of the MM-Fit poses. This difference likely hinders the model’s ability to learn when no calibration is performed. Indeed, calibration rectifies this issue. Excluding the scenarios without calibration for pose and multimodal inputs, our method surpasses the baseline in almost every case. The only exception is observed when training with sensor input and 100% training data, excluding the “No Activity” class and applying supervised calibration, where the baseline achieves a slightly better result (0.822 versus 0.798), which is due to the fact that MM-Fit uses real, rather than simulated sensor data.
Leveraging multimodal inputs (i.e., concatenating all modalities) achieves the highest performance, suggesting that HAR benefits from the diverse perspectives of multiple modalities. In this multimodal scenario, our method, compared to the baseline, improves the Macro -Score from 0.945 to 0.965 (with “No Activity”) and from 0.973 to 0.999 (without “No Activity”) when trained on the full dataset. Furthermore, with only 2% of the training data available, we improve performance from 0.744 to 0.932 (with “No Activity”) and from 0.819 to 0.992 (without “No Activity”). These findings show that our model has significantly better performance compared to the baseline and learns more efficiently when only a small amount of training data is available.
Figure 4 illustrates boxplots that exemplify the classification performances on accelerometer data input for all training fractions. It shows that our model significantly outperforms the baseline when trained with a minimal amount of training data, and generally, we observe a notable reduction in variance for our pretrained models compared to the baseline. This suggests that our pretrained sensor encoder has effectively learned a meaningful representation of real accelerometer data.
Table 6.2 demonstrates the generalization performance of our representation on MM-Fit in the scenario where the dataset is unseen. In this analysis, we contrast the outcomes of unimodal training—where models were individually trained to align sensor-text, pose-text, and video-text—with those of multimodal training, where we tasked a single model to align all four modalities concurrently. Our findings reveal that, after implementing supervised calibration for pose and sensor modalities, our model surpasses random chance performance in all but one scenario. The peak top-1 performance achieves a Macro -Score of 0.638 for pose input. The only exception is the accelerometer input when the “No Activity” category is included. In this case, the top-1 performance is for unimodal and for multimodal training, both of wich are less than . This is also the only instance where multimodal training does not yield a performance enhancement, being outperformed by a small margin of 0.005. Conversely, in all other training configurations, multimodal training demonstrates superior performance over unimodal training. On average, multimodal training increases the performance by 0.069 (top-1). This supports our narrative that multimodal training can augment the model’s comprehension of various inputs.
|c|c|c|ccc|ccc|
Input Modality Calibration Macro -Score
Unimodal Training Multimodal Training
Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
With “No Act.”
Pose No
Unsupervised
Supervised
Video No
Sensor No
Unsupervised
Supervised
Without “No Act.”
Pose No
Unsupervised
Supervised
Video No
Sensor No
Unsupervised
Supervised
7 Conclusion
In this study, we have developed a novel approach to HAR by introducing a joint feature space that integrates video, pose, simulated accelerometer, and textual data. Our model was trained on a large, self-created multimodal fitness dataset containing over 78 hours of recordings. We evaluated our method using the FLAG3D dataset and the MM-Fit dataset. On MM-Fit, our model achieved an impressive Macro -Score of up to 0.992 with only 2% of the train data and 0.999 when all available training data was used. Moreover, in the scenario where the MM-Fit dataset was unseen, we demonstrated a generalization performance of up to 0.638. Our results reveal that our pretrained encoders effectively capture a meaningful representation for recognizing human activities. Furthermore, our approach not only demonstrates impressive results on an unseen dataset but also is capable of enhancing performance when applied to real-world sensor data.

### v2

@@
 Abstract
-Human Activity Recognition is a longstanding problem in AI with applications in a broad range of areas: from healthcare, sports and fitness, security, and human computer interaction to robotics.
+Human Activity Recognition (HAR) is a longstanding problem in AI with applications in a broad range of areas, including healthcare, sports and fitness, security, and more.
 The performance of HAR in real-world settings is strongly dependent on the type and quality of the input signal that can be acquired.
-Given an unobstructed, high-quality camera view of a scene, computer vision systems, in particular in conjunction with foundational models (e.g., CLIP), can today fairly reliably distinguish complex activities.
-On the other hand, recognition using modalities such as wearable sensors (which are often more broadly available, e.g, in mobile phones and smartwatches) is a more difficult problem, as the signals often contain less information and labeled training data is more difficult to acquire.
-In this work, we show how we can improve HAR performance across different modalities using multimodal contrastive pretraining.
-Our approach MuJo (Multimodal Joint Feature Space Learning), learns a multimodal joint feature space with video, language, pose, and IMU sensor data.
-The proposed approach combines contrastive and multitask learning methods and analyzes different multitasking strategies for learning a compact shared representation.
-A large dataset with parallel video, language, pose, and sensor data points is also introduced to support the research, along with an analysis of the robustness of the multimodal joint space for modal-incomplete and low-resource data.
-On the MM-Fit dataset, our model achieves an impressive Macro -Score of up to 0.992 with only 2% of the train data and 0.999 when using all available training data for classification tasks.
-Moreover, in the scenario where the MM-Fit dataset is unseen, we demonstrate a generalization performance of up to 0.638.
+Given an unobstructed, high-quality camera view of a scene, computer vision systems, in particular in conjunction with foundation models, can today fairly reliably distinguish complex activities.
+On the other hand, recognition using modalities such as wearable sensors (which are often more broadly available, e.g., in mobile phones and smartwatches) is a more difficult problem, as the signals often contain less information and labeled training data is more difficult to acquire.
+To alleviate the need for labeled data, we introduce our comprehensive Fitness Multimodal Activity Dataset (FiMAD) in this work, which can be used with the proposed pre-training method MuJo (Multimodal Joint Feature Space Learning) to enhance HAR performance across various modalities.
+FiMAD was created using YouTube fitness videos and contains parallel video, language, pose, and simulated IMU sensor data.
+MuJo utilizes this dataset to learn a joint feature space for these modalities.
+We show that classifiers pre-trained on FiMAD can increase the performance on real HAR datasets such as MM-Fit, MyoGym, MotionSense, and MHEALTH.
+For instance, on MM-Fit, we achieve an Macro -Score of up to 0.855 when fine-tuning on only 2% of the training data and 0.942 when utilizing the full training set for classification tasks.
+We have compared our approach to other self-supervised ones and showed that, unlike them, ours can consistently improve on the baseline network performance as well as provide a better data-efficiency.
 1 Introduction
-Perceiving and interpreting human activity is a core functionality of intelligent systems with applications in a broad range of areas: from healthcare Tan et al.
-(2021) through sports and fitness Host and Ivašić-Kos (2022); Nadeem et al.
-(2020), security Sunil et al.
-(2021) to robotics Piyathilaka and Kodagoda (2015).
-A basic capability needed to facilitate such functionality is the identification of concrete human actions from the system’s sensory input, referred to as Human Activity Recognition (HAR) Hamad et al.
-(2023).
+Perceiving and interpreting human activities is a core functionality of intelligent systems with applications in a broad range of areas, including healthcare Tan et al.
+(2021), sports and fitness Host and Ivašić-Kos (2022); Nadeem et al.
+(2020), and security Sunil et al.
+(2021), among others.
+A basic capability needed to facilitate such functionality is the identification of concrete human actions from the system’s sensory input, referred to as Human Activity Recognition (HAR).
 In general, HAR is a difficult problem, especially when applied to uncontrolled real-life environments, where it must contend with the variability of human actions and the open-ended variations in environmental conditions.
 In such settings, the performance of HAR systems is strongly dependent on the available input modalities.
-Recent advances in Computer Vision, particularly with Large Vision Language Models (VLMs) Radford et al.
+Recent advances in Computer Vision, particularly with Large Vision-Language Models (LVLMs) Radford et al.
 (2021); Achiam et al.
@@
 Much more broadly available input modalities are wearable sensors, in particular, Inertial Measurement Units (IMUs), which are ubiquitous in smartphones, smartwatches or even earbuds, all of which people carry with them on a daily basis.
-Unfortunately, achieving acceptable recognition performance with such sensors is a much more difficult problem due to two factors:
-1.
-Information ambiguity: An IMU (which is essentially a motion sensor) in a smartwatch on the wrist or in a smartphone in a pocket contains much less detailed information about the user’s activity than a high-quality picture.
-2.
-Lack of training data: The success of VLMs is built on the online availability of massive amounts of text/image data pairs, which can be used for training aligned representations.
+Unfortunately, achieving acceptable recognition performance with such sensors is a much more challenging problem due to two factors, the first one being information ambiguity.
+An IMU (a motion sensor) in a smartwatch on the wrist or in a smartphone in a pocket contains much less detailed information about the user’s activity than a high-quality picture.
+This problem is heightened by the second factor, lack of training data.
+The success of LVLMs is built on the online availability of massive amounts of text/image data pairs, which can be used for training aligned representations.
 Although large amounts of unlabeled IMU data are increasingly becoming available Chan et al.
 (2024), labeled training data remains scarce.
-The work presented in this paper, which aims to address the issues outlined previously, is founded on three key observations:
+The work presented in this paper, which aims to address these issues, is founded on three key observations:
 1.
-While the input modalities available during inference—the time when the system is utilized in the real world—may be constrained, the input modalities during training can be selected with much greater flexibility.
+While the input modalities availability during inference—the time when the system is utilized in the real world—may be constrained, the input modalities during training can be selected with much greater flexibility.
 2.
@@
 (2019) are used to generate skeletal data of human activities.
-This enables the creation of synchronised video/sensor training data and through the alignment of video with text, it facilitates the synchronization of sensor data with textual descriptions.
-Based on the above considerations, we propose a multimodal contrastive pretraining method that creates a joint representation between videos, video-derived poses, synthetic IMU data created from videos using IMUTube Kwon et al.
-(2021), and textual descriptions.
-Our approach, which we refer to as MuJo, aims to utilize the complementary information of each modality to improve the performance for both unimodal and multimodal inference.
+This enables the creation of synchronized video/sensor training data and through the alignment of video with text, it facilitates the synchronization of sensor data with textual descriptions.
+Based on the above considerations, we propose a Multimodal Joint Feature Space Learning method (MuJo) that learns a joint representation between videos, video-derived poses, virtual IMU data generated from videos, and textual descriptions by employing pair-wise contrastive learning, as depicted in Figure 1.
+MuJo aims to utilize the complementary information of each modality to improve the performance for both unimodal and multimodal inference.
 In particular, we show that our method dramatically reduces the need for labeled training data for downstream tasks.
-Thus, on the well-known MM-Fit dataset Strömbäck et al.
-(2020), when using all available training data, our method improves the recognition rate (Macro -Score) from 0.973 to 0.999 in the multimodal case (encompassing all of the above modalities), from 0.957 to 0.988 for poses, and from 0.727 to 0.863 for video alone.
-For sensor data, the performance of our method is slightly worse (0.798 as opposed to 0.822), which is due to the fact that MM-Fit uses real, rather than simulated sensor data.
-However, when using just 2% of the available data for training, the baseline with sensors has an -Score of just 0.378 whereas our systems achieves a reasonable 0.627 (see section 6.2 for details).
-The multimodal -Score is 0.819 in the baseline, while our method raises it to 0.992.
+For that, we provide extensive evaluation on four publicly available HAR datasets, namely MM-Fit Strömbäck et al.
+(2020), MyoGym Koskimäki et al.
+(2017), MotionSense Malekzadeh et al.
+(2018), and MHEALTH Banos et al.
+(2015).
+For instance, on the MM-Fit dataset in the multimodal case (encompassing all of the aforementioned modalities as input), our method improves the recognition rate (Macro -Score) from 0.897 to 0.942 when using all available training data.
+Even when utilizing only 2% of the training data, it enhances the recognition rate from 0.790 to 0.855.
 Overall, the key contributions of this work can be summarized as follows:
-Development of a large fitness dataset111Available after publication at link.
-with parallel video, language (i.e., labels of video clips, captions uttered by the instructor in the video clip, and precise descriptions of the exercises generated by GPT-3.5), pose, and simulated IMU data points.
-Learning of an effective joint space for the modalities of text, video, pose, and sensor by leveraging the contrastive loss.
-Analysis of usefulness and generalizability of multimodal joint space for each modality and their combination, as well as the benefits of pretraining when a low amount of data is available for the target task.
+1.
+Development of the Fitness Multimodal Activity Dataset (FiMAD)111Available after publication at link., a comprehensive dataset containing parallel video, language (including video clip labels, instructor utterances, and detailed exercise descriptions generated by GPT-3.5 Turbo), pose data, and simulated IMU data points.
+2.
+Pe-training of MuJo, a multimodal joint space model, using the FiMAD dataset.
+MuJo integrates four different modalities: text, video, pose, and virtual sensor data, leveraging contrastive learning inspired by CLIP Radford et al.
+(2021).
+3.
+Evaluation of the utility and generalizability of MuJo for real-world accelerometer data on the MM-Fit, MyoGym, MotionSense, and MHEALTH datasets.
+On the MM-Fit dataset, we additionally conducted experiments with pose, video, and multimodal inputs.
+The benefits of pre-training MuJo on FiMAD, especially when limited data is available for the target task, are analyzed.
+Furthermore, we analyze the zero-shot performance of our method and compare unimodal versus multimodal training on all four datasets.
 2 Related Work
-2.1 Human Activity Recognition
-Unimodal activity recognition.
+Unimodal and multimodal activity recognition.
 Unimodal activity recognition has been extensively studied in HAR.
@@
 CNNs are commonly used for image-based activity recognition Strömbäck et al.
-(2020); Ordóñez and Roggen (2016), being combined with RNNs in some cases Ordóñez and Roggen (2016).
+(2020), being combined with RNNs in some cases Ordóñez and Roggen (2016).
 Despite the success of unimodal activity recognition, it is often limited by the fact that a single modality of input data may only capture some of the necessary information about the activity, leading to lower recognition accuracy.
 Therefore, multimodal activity recognition, which combines information from multiple sources, has gained increasing attention in recent years.
-Multimodal activity recognition.
-Multimodal activity recognition recognizes human activities using information from multiple modalities, such as IMU data, audio, and video Mekruksavanich and Jitpattanakul (2022); Ijaz et al.
+Multimodal activity recognition, on the other hand, recognizes human activities using information from multiple modalities, such as IMU data, audio, and video Mekruksavanich and Jitpattanakul (2022); Ijaz et al.
 (2022).
@@
 (2022).
-2.2 Representation Learning for Human Activities
 Representation learning with proxy tasks.
@@
 Generating one feature vector with multimodal inputs for a common concept is a crucial problem of multimodal representation learning.
-Simple methods like concatenating Hu et al.
+Simple methods, such as concatenating learned uni-modal feature vectors, have been extensively explored to enhance the performance of downstream tasks Hu et al.
 (2018); Strömbäck et al.
-(2020) learned uni-modal feature vectors have been extensively explored to enhance the performance of downstream tasks.
+(2020).
 These methods involve combining feature vectors from a single modality, such as text or image, to create more informative representations that can be used for various tasks.
-Thereafter, Zadeh et al.
-proposed a tensor fusion network in Zadeh et al.
+More advanced methods have also been proposed, including a tensor fusion network Zadeh et al.
+(2017) and a transformer-based co-attention mechanism for efficient cross-modality learning Lu et al.
 (2019).
-Lu et al.
-proposed a co-attention mechanism based on the transformer architecture in Lu et al.
-(2019) for efficient cross-modality learning.
+Proxy task were also developed for the field of Human Activity Recognition with weareable devices including IMUs, where unlabelled IMU data is used to pre-train better representations.
+Those approaches include Multitask Saeed et al.
+(2019), Auto-encoder Haresamudram et al.
+(2019) and SimClr Tang et al.
+(2021).
+Multi-task Saeed et al.
+(2019) proposes auxiliary tasks of binary classification for extracting useful features for the downstream task.
+A multi-task temporal convolutional neural network (CNN) is trained to classify whether the input signal was transformed or not for each of a set of transformations such as scaling, rotating the signal, etc.
+The authors utilize a large unlabelled accelerometer dataset for self-supervised learning and transfer learned knowledge to a small, labelled activity recognition dataset.
+Auto-encoder Haresamudram et al.
+(2019) is the application of auto-encoders where an encoder network reduces the dimensionality of the data while a decoder one brings the representation back to its original dimensions, with the whole system learning to reconstruct the original signal.
+SimClr Tang et al.
+(2021) learns general representations from a dataset where a model is trained to match different views instances created from the same signal by contrasting them.
+The positive pair is obtained by the transformation of the input window, while the negative pair is formed by the rest input windows of the same batch.
+Then, the contrastive loss pulls the positive pair together while pulling the negative pair far away from each other in the representation space.
 Representation via contrastive learning.
-Contrastive Learning is as an efficient self-supervised framework applied across multiple domains, which learns similar/dissimilar representations from data that are organized into similar/dissimilar pairs.
+Contrastive learning is as an efficient self-supervised framework applied across multiple domains, which learns similar/dissimilar representations from data that are organized into similar and dissimilar pairs.
 Recently, Radford et al.
@@
 CLIP learns two feature spaces with two encoders, i.e., image and text encoders, then projects the learned features into a shared latent space with contrastive loss.
-Contrastive learning, a joint feature space of text-image pairs, has been extended to other modalities.
-For example, video-text data has attracted studies on shared space of video-text features Xu et al.
+Contrastive learning has been extended to other modalities as well.
+For example, video-text data has attracted interest in studies focused on the shared space of video-text features Xu et al.
 (2021); Xue et al.
-(2022), text-to-video generation Singer et al.
+(2022) and text-to-video generation Singer et al.
 (2022).
-Thereafter, MotionCLIP Tevet et al.
-(2022) trained an encoder to find the proper embedding of an input sequence in CLIP space, and a decoder that generates the most fitting motion to a given CLIP space.
-Following that, IMU2CLIP Moon et al.
-(2022) has been proposed to learn a joint space to align Inertial Measurement Unit (IMU) motion sensor recordings with video and text, by projecting them into the joint representation space.
-Studies in action recognition Aggarwal and Ryoo (2011); Li et al.
+MotionCLIP Tevet et al.
+(2022) trained an encoder to find the appropriate embedding of an input sequence in CLIP space and a decoder to generate the most fitting motion for a given CLIP space.
+Following this, IMU2CLIP Moon et al.
+(2022) was proposed to learn a joint space that aligns IMU motion sensor recordings with video and text by projecting them into a shared representation space.
+Studies in action recognition Li et al.
 (2020); Wang et al.
@@
 (2022); Morshed et al.
-(2023) can also significantly enhance the field of recognition in fitness activities, despite action recognition emphasis on a broad spectrum of actions such as “cutting the potato” and “cleaning the drawer”, among others.
-Certain investigations in action recognition also endeavor to acquire knowledge of the feature space through contrastive learning methodologies, exemplified by initiatives such as HERO Li et al.
+(2023) have also significantly enhanced the field of fitness activity recognition, despite traditionally focusing on a broad spectrum of actions such as “cutting the potato” and “cleaning the drawer”.
+Certain investigations in action recognition also endeavor to acquire knowledge of the feature space through contrastive learning, as exemplified by initiatives like HERO Li et al.
 (2020) and ActionCLIP Wang et al.
 (2021).
-The achievements discussed in the aforementioned studies have shown contrastive learning has a high impact on extracting a joint space with similar and dissimilar pairs.
-Hence, the contrastive learning methodology employed in this study is utilized to acquire knowledge of the feature space pertaining to fitness activities.
+The achievements discussed in the aforementioned studies have shown that contrastive learning has a significant impact on extracting a joint space with similar and dissimilar pairs.
+Consequently, the contrastive learning methodology employed in our study aims to acquire knowledge of the feature space pre-training to fitness activities.
 Contrary to previous work, we extend this line of work on contrastive learning to a unique multimodal setting that utilizes more than two modalities, such as video, IMU sensors, poses, and text, within a multitask training manner in an end-to-end training paradigm similar to ImageBind Girdhar et al.
@@
 Thus, our model and data provide a cross-modal classification opportunity.
-3 Dataset
-3.1 Video Collection
+3 Dataset Collection and Construction
+Annotating video descriptions is time-consuming and expensive.
+Thus, we create our dataset from a comprehensive and varied collection of videos with temporally localized chapter information, including chapter titles and start times, specifically for fitness activities.
 YouTube provides chapters within a video, each with its preview, providing information and context for different parts of the video.
-Video owner can manually add chapters to their uploaded videos or rely on automatic chapter generation.
-We manually collect 1212 fitness videos with pre-annotated video chapters.
+Those chapters are manually added by the uploader or automatically generated.
+Each chapter is a continuous, non-overlapping segment that divides the video into a video clip and its corresponding textual description.
+We manually collect 1388 fitness videos with pre-annotated video chapters.
 Additionally, we retrieve the captions automatically generated by YouTube.
-Our videos include only one tutor who teaches how to do fitness exercises (e.g., “squat with dumbbells”, “sit-ups”, “side deep squats”, etc.).
-Thus, the exercise videos provide instructions for the activities while showing the activities clearly.
+All our videos show a tutor who teaches how to do fitness exercises (e.g., “squat with dumbbells”, “sit-ups”, “side deep squats”, etc.).
+The trainers are primarily performing HIIT home workouts, focusing on bodyweight exercises with no or less equipment.
+The exercise videos provide instructions for the activities while showing the activities clearly.
 The annotation does not include predefined categories or labels.
 Each video owner annotates the activity segment of the videos with their personal preferences, such as “deep squats with dumbbells”, “high jump squats”, etc., instead of only “squats”.
-Given the temporal boundaries of the activity within each video, we split the videos in shorter activity clips along with their corresponding chapters and caption segments.
-Since workout exercises are typically short, activity videos that are too long often contain multiple exercises and, thus, incorrect chapter labels.
-Hence, we only keep exercises shorter than 2 minutes and discard more extended exercises.
-After that, the dataset consists of 10,695 (video, chapter, caption) samples.
-3.2 Dataset Construction
-We convert all videos to 720p resolution and 50 frames per second via FFmpegs adaptive overlapped block motion compensation with bidirectional motion estimation.
-Then, we extract simulated sensor data, poses, as well as a variety of video and text features.
-Video features are extracted with HERO Li et al.
-(2020) utilizing CLIP Radford et al.
-(2021) and S3D Xie et al.
+Based on the temporal boundaries of the chapters within each video, we split the videos in shorter activity clips along with their corresponding caption segments.
+Since workout exercises are typically short, activity videos that are too long often contain multiple exercises and, thus, incorrect chapter annotations.
+Hence, we only keep exercises shorter than 2 minutes and discard longer exercises.
+After that, the dataset consists of 10,695 (video, chapter, caption)-samples.
+We convert all 10,695 exercise videos to 720p resolution and 50 frames per second via FFmpegs adaptive overlapped block motion compensation with bidirectional motion estimation.
+To enhance the dataset with diverse modalities, we use IMUTube Kwon et al.
+(2021) to obtain pose and virtual sensor (accelerometer and gyroscope) features.
+As a result, our dataset includes (video, pose, sensor, text)-pairs for each input.
+For video and text we extracte multiple features.
+Video features include using HERO Li et al.
+(2020), which is utilizing either CLIP Radford et al.
+(2021) or S3D Xie et al.
 (2018) embeddings.
-On the textual side, we apply several normalization strategies to the raw chapters such as replacing abbreviations within the text (e.g., changing “l” to “left” and “r” to “right”) as well as removing stop words, numbers, and a few special symbols from a predefined list.
-Further, we use GPT 3.5 to generate detailed descriptions of the exercises.
-We then extract multiple text features with different models, including OpenAI’s text-embedding-ada-002 embedding model (ADA), CLIP, BERT Devlin et al.
-(2018), as well as language features and motion tokens extracted with Text2Motion Transformer Guo et al.
+For text, we apply several normalization strategies to the raw chapters such as replacing abbreviations within the text (e.g., changing “l” to “left” and “r” to “right”) as well as removing stop words, numbers, and a few special symbols from a predefined list.
+Further, we use GPT-3.5 Turbo to generate detailed descriptions of the exercises.
+We then extract multiple text features with various models, including OpenAI’s text-embedding-ada-002 embedding model (ADA), CLIP, BERT, as well as language features and motion tokens extracted with the Text2Motion Transformer Guo et al.
 (2022).
-We apply IMUTube Kwon et al.
-(2021) to generate virtual gyroscope and accelerometer data for sensors placed on all 16 joint nodes.
+We apply IMUTube to generate virtual gyroscope and accelerometer data for sensors placed on all 16 joint nodes.
 We also use the implementation of IMUTube to extract 3D poses from the videos, which utilizes OpenPose followed by VideoPose3D for 3D pose lifting.
-Since IMUTube fails to extract sensor data for 884 videos, we exclude these sequences.
-We use the remaining 9811 exercises as our dataset and split it into train (8,518 videos), validation (946 videos) and test (347 videos) sets.
-Since workout session videos typically start with an introductory sequence and conclude with a farewell, we discard the initial and final 15%222This value has proven to be effective in experimental investigations.
-of each video.
+Since IMUTube fails to extract sensor data for 884 exercise videos, we exclude these sequences.
+We use the remaining 9811 exercises as our dataset and split it into 8,227 train and 1,584 validation activity clips.
+The split is based on the level of the 1,388 original fitness videos, to ensure video independence.
+We do not create a test set because the dataset is used solely for pre-training, not for testing.
+Since workout session videos typically start and end with explanations given by the trainer, we discard the initial and final 15%222This value has proven to be effective in experimental investigations.
+of each activity clip.
 Despite these filtering steps, we receive a large dataset totaling around 78 hours in length.
@@
 We found this size suitable for fitness exercises as a 2 second window generally captures enough meaningful information regarding the subject’s movements.
-In summary, our dataset comprises a large collection of short fitness activity clips containing video features, poses, synthetic IMU data, chapter labels, captions, and GPT-generated descriptions as well as several types of features extracted from these text modalities.
+In summary, our dataset comprises a large collection of short fitness activity clips containing video features, poses, virtual IMU data, chapter labels, captions, and GPT-generated descriptions as well as several types of features extracted from these text modalities.
 4 Method
-Our method leverages multimodal information for HAR pretraining.
-The fundamental idea is that the extracted information for a short video interval should be similar, as it represents the same activity from various perspectives.
-For each window, we extract features for each modality, which means our data consists of where is the feature for modality at time interval .
-We use as modalities
-: The sequence of 3D poses for the subject extracted using VideoPose3D.
-: Simulated accelerometer data of both wrists generated by ImuTube Kwon et al. (2021).
-: Embeddings extracted using OpenAI’s ADA model from YouTube’s chapter information with a lightweight normalization process removing a few special symbols and replacing abbreviations within the text.
+Our method leverages multimodal information for HAR pre-training.
+In this Section, we outline the task and its formulation, along with detailing the multimodal representation methods employing CLIP Radford et al.
+(2021)-like contrastive learning.
+Our objective is to improve HAR by employing contrastive learning across multiple modalities.
+The fundamental idea is that the extracted information from different modalities for a short video interval should be similar, as they represent the same activity from various perspectives.
+Each instance is a window, and for each one, we extract features for every modality.
+This means our data consists of , where is the feature for modality at window .
+We use as modalities:
+: Simulated accelerometer data generated by ImuTube.
+: Embeddings extracted using OpenAI’s ADA model from YouTube’s chapter information with a lightweight normalization process removing a few special symbols and replacing abbreviations.
 Each modality has its own encoder network and projection layers .
 Thus, serves as our representation for a vector of modality in our shared representation space.
-We learn this shared space between modality representations using pair-wise contrastive learning, as depicted in Figure 1.
+We learn this shared space between modality representations using the InfoNCE loss.
 First, for all possible modality pairs, we compute similarities using
@@
 For and , we employ the CNN encoder architecture as described in Strömbäck et al.
-(2020).
+(2020), making slight adaptations to accommodate our input shapes.
 The projection is a two-layer fully connected model for all modalities , mapping them to feature vectors of length 1280.
-We apply a random search approach for hyperparameter selection.
+Hyperparameter and model optimization.
+We applied a random search approach for hyperparameter selection.
 Our hyperparameters include learning rate, number and size of projection layers, dropout rates, and activation functions.
-We also experimented with various configurations of the loss function, specifically comparing using the text modality as an anchor and, hence, considering sensor-text, pose-text, and video-text loss pairs only versus incorporating all possible modality combinations into the loss.
+Furthermore, we experimented with various configurations of the loss function, specifically comparing the effect of using the text modality as an anchor, i.e., only considering sensor-text, pose-text, and video-text loss pairs, versus incorporating all possible modality combinations into the loss.
 Encompassing all modality pairs together yielded superior results.
 We explored weighting the text modality higher than the others in the loss, but it did not have any observable effect on performance.
-We also explored the use of S3D-based video embeddings, the gyroscope as sensor input, and the other text modalities and features described in Section 3.2 as inputs.
-However, since those configurations did not yield the best performance during the hyper-parameter search, we have omitted those results for the sake of brevity.
+We also explored the use of S3D-based video embeddings, the gyroscope as sensor input, and the other text modalities and features described in Section 3 as inputs.
+However, since those configurations did not yield optimal performance during the hyper-parameter search, we have omitted those results for the sake of brevity.
+5 Evaluation Methodology
+5.1 Evaluation Datasets
+We conduct experiments on the MM-Fit Strömbäck et al.
+(2020), MyoGym Koskimäki et al.
+(2017), MotionSense Malekzadeh et al.
+(2018), and MHEALTH Banos et al.
+(2015) datasets.
+MM-Fit is designed for multimodal fitness activity recognition and combines data from multiple modalities, including video, pose estimations, and various sensors.
+The dataset includes 21 workout sessions conducted by ten subjects, each performing ten different fitness exercises.
+The MyoGym dataset comprises 30 distinct gym exercises focused on upper body training by ten individuals, each equipped with a sensor on their right forearm.
+The MotionSense dataset consists of sensor data collected from 24 participants carrying an iPhone 6s in their front pocket, covering six different activities (walking downstairs, walking upstairs, sitting, standing, and jogging).
+The MHEALTH dataset includes twelve physical activities (such as climbing stairs, cycling, and running) performed by ten subjects, with data captured by three sensors positioned on the chest, right wrist, and left ankle.
+All datasets, except for MotionSense, include a "No Activity" category (NULL class) in addition to the specified activities.
+5.2 Unimodal Classification on Real-World Accelerometer Data
+We evaluate the performance of our pre-trained representation on real-world accelerometer data using all four datasets.
+For MM-Fit, we use the accelerometer data from the smartwatches worn on the left and right wrists, which we had to downsample from 100 Hz to 50 Hz via linear interpolation to match our input frequency.
+For MyoGym, MotionSense, and MHEALTH, we use accelerometer data from all available sensors, which are already at 50 Hz, allowing us to employ the data directly without additional preprocessing.
+For all four datasets, we apply -score normalization and use a sliding window of 100 samples in length (i.e., 2 seconds) with 50 samples overlap to generate instances.
+Furthermore, since each target dataset has different sensor placements, we have to pre-train a separate MuJo model for each one.
+Specifically, during pre-training our models, we use the virtual sensor positions generated by IMUTube that best match the sensor positions of the respective target dataset.
+Each of the four models is pre-trained on our FiMAD dataset incorporating all four input modalities: text, video, poses, and virtual accelerometer data.
+We conduct our experiments based on the standard experimental methodology in the HAR community, i.e., testing on unseen users and evaluating using the Macro -Score, a metric that is well-suited for handling class imbalance.
+For MM-Fit, we use the training, validation, and cross-subject test sets as described in Strömbäck et al.
+(2020).
+Since no fixed splits are provided for MyoGym, MotionSense, and MHEALTH, we employ -fold cross-validation with .
+Additionally, we consider two scenarios: training the classifiers with and without the inclusion of the NULL class.
+This is relevant because the NULL class is not part of our pre-training dataset but is present in real-world scenarios and included in all evaluation datasets except for MotionSense.
+For classification, we utilize a classifier architecture comprising our sensor encoder , projection , and a classification head consisting of 2 layers.
+We contrast three scenarios: In the baseline setting, all weights are initialized randomly, meaning the model lacks any pre-trained knowledge.
+In the other two scenarios, the encoder and projection head are initialized with weights from our pre-trained model.
+In one case, the weights of the encoder and projection head are fixed (frozen), whereas in the other case, they are trainable and we fine-tune them during the classifier’s training process.
+We train with different proportions of the training data, ranging from one to 100 percent, with the training data being sampled randomly per class to ensure balanced representation.
+To guarantee comparability of all results, we train each classifier with the same settings: Training is conducted for a maximum of 200 epochs, incorporating early stopping after 25 epochs without improvement.
+We employ a weighted cross-entropy classification loss to address class imbalance, given by:
+where is the number of classes, is a one-hot encoded ground truth activity label and is the predicted label vector of for input , that is, the application of our classification layers to obtain .
+The weight for class is calculated by dividing the number of data samples of the largest class by the number of instances in class itself.
+To ensure validity of the results, we train 20 models independently for each configuration, with the training data being randomly selected anew for each run (for proportions less than 100%) and the non-pre-trained weights being reinitialized randomly.
+5.3 Proxy task comparisons
+We compare our approach with Multi-task Saeed et al.
+(2019), Auto-encoder Haresamudram et al.
+(2019) and SimClr Tang et al.
+(2021).
+All approaches were trained with the same simulated IMU data as our approach and use the same encoder architecture as ours, for fair comparison.
+Regarding other configurations SimClr used random rotations as augmentation uniformely sampled in range.
+The representation is obtained from our projection layer through two dense layers of size 256, 128 and 128, respectively, with the first two being followed by a relu activation.
+It was trained with the NTXent loss using l2 norm.
+For Multitask we have selected as augmentations noising, scaling, negation, horizontally flipping, permutation and channel shuffling.
+Noising consists of adding a value sampled from a normal distribution with mean zero and standard deviation of 0.05 while scaling consists of multiplying by values drawn from this same distribution.
+Negation, horizontally flipping and channel shuffling are self evident.
+Permutation means randomly splitting the time series into four parts and changing their order.
+Every task has a head that takes as input the encoder’s output and consists of one fully connected layer with 256 size and relu activation, followed by another fully connected layer with also 256 size but sigmoid activation.
+For Auto-encoder the representation is passed to a convolutional decoder that re-creates the original signal.
+This decoder consists of first a linear layer to bring the projection to where is the number of sensors in the dataset and is the original window size.
+This is followed by batch norm, relu and then three blocks of convolution transposed, dropout, relu and upsampling.
+A kernel size of was used and the channels change following the sequence , with being the original number of sensor channels.
+To our knowledge, this is the first application of such methods to simulated sensor data, as well as their first use in a multi-sensor scenario.
+While all those methods could be added to our framework easily, we selected them as comparison to highlight how our method can learn from extra modalities only available at training time beyond what is present at the simulated sensor data itself (which is the only input for those methods).
+5.4 Multimodal Classification on MM-Fit
+We evaluate the multimodal capabilities of our system using the MM-Fit dataset.
+Since MM-Fit comes along with videos and poses at 30 Hz, we upsample both modalities to 50 Hz.
+We extract HERO CLIP features from the videos following the approach outlined in Section 3.
+We use the sliding window method described in Section 5.2 to generate instances for pose and video inputs.
+We also employ the same classifier architecture but select the encoders and projections depending on the input modality.
+Specifically, we train individual classifier models for the pose and video input modality, using their respective encoders and projections .
+Additionally, we experiment with multimodal input by providing all three modalities (sensor, pose, and video) to the model and concatenating the outputs from the respective projections before feeding them into the classifier head .
+We keep the same training setup as outlined for unimodal training in Section 5.2, perform training runs with and without the NULL class, and contrast the same three training scenarios: baseline, pre-trained, and pre-trained with frozen encoder and projection weights.
+For the pre-trained scenarios, we reuse the MuJo model from Section 5.2 that was pre-trained for the MM-Fit target dataset (i.e., allowing left and right wrist sensor placements).
+|c|c|c|ccc|ccc|
+Dataset
+With
+“NULL”
+Method Mean Macro -Score
+Train Data: Train Data:
+Baseline (
+) Pre-trained (
+) Pre-trained (
+) Baseline (
+) Pre-trained (
+) Pre-trained (
+MYOGYM ✗ MuJo
+Auto-encoder
+SimClr
+Multitask
+2-9
+✓ MuJo
+Auto-encoder
+SimClr
+Multitask
+MHEALTH ✗ MuJo
+Auto-encoder
+SimClr
+Multitask
+2-9
+✓ MuJo
+Auto-encoder
+SimClr
+Multitask
+MM-Fit ✗ MuJo
+Auto-encoder
+SimClr
+Multitask
+2-9
+✓ MuJo
+Auto-encoder
+SimClr
+Multitask
+MOTIONSENSE ✗ MuJo
+Auto-encoder
+SimClr
+Multitask
+5.5 Zero-Shot
+Beyond classification, we evaluate the ability of our pre-trained MuJo model to generalize to unseen datasets without additional fine-tuning.
+We first calculate the ADA text embeddings for all labels .
+For the NULL class, which doesn’t correspond to a specific activity, we generate its embedding by taking the mean of the embeddings of all labels.
+For a given data sample of input modality , we then choose labels based on the closest similarity score by calculating for all labels .
+We contrast the outcomes of unimodal training—where MuJo models were individually trained to align sensor-text, pose-text, or video-text—with those of multimodal training, where we tasked a single model to align all four modalities concurrently.
+We conduct experiments on all four datasets, both with and without the NULL class, and calculate the accuracy and Macro -Score.
+6 Results and Discussion
+6.1 Unimodal Classification on Real-World Accelerometer Data
+Table 5.4 shows that for our method the pre-trained models with trainable weights consistently outperform the baseline with respect to Macro -Score on all dataset, regardless of whether 2% or 100% of the training data is used.
+These findings indicate that multimodal pre-training via contrastive learning on virtually generated accelerometer data can effectively improve HAR performance on real-world sensor data.
+Regarding comparison with other approaches, we can see how our method is more data-efficient, as we outperform other approaches in almost all cases when 2% of the labeled training data is available.
+When 100% of the training data is available every method is dominant in a different dataset, but ours consistently improves on the baseline unlike other approaches.
+Moreover, our method presents a more consistent performance when freezing the encoder, showing the quality of the learned representation.
+In general including the NULL class generally results in lower scores, but our trainable models still achieve higher scores than the baseline, even if this class was not part of our pre-training.
+Only for the MHEALTH dataset, when utilizing 100% of the data, the performance difference becomes marginal, with scores of 0.839 versus 0.837 without the NULL class and 0.582 versus 0.577 with the NULL class.
+However, when labeled data is limited, we perform significantly better than the baseline.
+For example, when utilizing only 2% of the train data, even the pre-trained classifiers with frozen encoder and projection layers outperform the baseline in all cases except for the MM-Fit dataset when the NULL class is included.
+The boxplots presented in Figure 2 compare the baseline to classifiers pre-trained with MuJo, both with frozen and trainable encoders and projections.
+For various training fractions ranging from 1% to 100%, we observe that the performance of our frozen representations highly depend on the target dataset and whether the NULL class is included during training.
+For instance, when the NULL class is excluded, the frozen models perform comparable to the pre-trained trainable models on MYOGYM and only slightly worse on MM-Fit.
+On MYOGYM with the NULL class, the frozen classifiers perform comparably to the baseline, whereas in all other scenarios, they generally outperform the baseline when less training data is utilized, but perform worse when more data is included in the training.
+On the other hand, the pre-trained models with trainable weights constantly achieve higher scores than the baseline for all training fractions on all datasets, which demonstrates the effectiveness of our pre-trained representations for accelerometer-based activity recognition, and robustness to varying sensor placements and target domains.
+We have also compared MuJo to the other pre-trained methods with different training data fractions, all with trainable encoders, see Figure 3.
+Here it is clear that our method is clearly the best in the MYOGYM dataset.
+Moreover, it outperforms all others up to 10% training fraction in all dataset other than MM-Fit, where only Multitask is better.
+6.2 Multimodal Classification on MM-Fit
+Table 6.2 presents the results of our experiments conducted for pose, video, and multimodal (concatenation of sensor, pose, and video) inputs, using both 2% and 100% of the training data, and with and without the inclusion of the NULL class.
+While the classifiers with our frozen pre-trained MuJo encoders and projections often perform worse than the baseline, our trainable models outperform the baseline in all cases except for the pose input without the NULL class when using the full training data.
+In general, for the pose input, the performance difference between these two training scenarios is only marginal, as evident from Figure 4.
+This discrepancy may be attributed to differences between the poses generated by IMUTube and that extracted by MM-Fit.
+Although both IMUTube and MM-Fit utilize VidePose3D for pose extraction, IMUTube incorporates extensive additional pre- and post-processing steps.
+This hypothesis is supported by the results for the video modality, where our trainable models outperform the baseline in all scenarios.
+Since the video embeddings, which we used as inputs, were extracted in the same manner for both our FiMAD dataset and MM-Fit, there is no such inconsistency, unlike the poses.
+Leveraging multimodal inputs achieves the highest performance, suggesting that HAR benefits from the diverse perspectives of multiple modalities.
+In this multimodal scenario, our method, compared to the baseline, improves the Macro -Score from 0.824 to 0.882 (with NULL) and from 0.897 to 0.942 (without NULL) when trained on the full dataset.
+Furthermore, with only 2% of the training data available, we improve performance from 0.656 to 0.734 (with NULL) and from 0.790 to 0.885 (without NULL).
+These findings show that the pre-trained classifiers significantly outperform the baseline and learn more efficiently with a limited amount of training data across multiple input modalities.
 |c|c|ccc|ccc|
-Dataset Type Input Modality Mean Macro -Score
+With
+“NULL”
+Input Modality Mean Macro -Score
 Train Data: Train Data:
-Baseline Ours (Frozen) Ours Baseline Ours (Frozen) Ours
-Real-World Sensor
+Baseline (
+) Pre-trained (
+) Pre-trained (
+) Baseline (
+) Pre-trained (
+) Pre-trained (
 Pose
@@
 Multimodal
-Rendered Sensor
 Pose
@@
 Multimodal
-|c|c|c|ccc|ccc|
-Input Modality Calibration Mean Macro -Score
-Train Data: Train Data:
-Baseline Ours (Frozen) Ours Baseline Ours (Frozen) Ours
-Without “No Activity”
-Sensor No
-Unsupervised
-Supervised
-Pose No
-Unsupervised
-Supervised
-Video No
-Multimodal No
-Unsupervised
-Supervised
-With “No Activity”
-Sensor No
-Unsupervised
-Supervised
-Pose No
-Unsupervised
-Supervised
-Video No
-Multimodal No
-Unsupervised
-Supervised
-5 Experiments
-We evaluate the performance of our methodology on the FLAG3D and MM-Fit datasets.
-On FLAG3D, we show performance in classification task whereas on MM-Fit, besides classification, we additionally evaluate the generalization performance of our representation in the scenario where the dataset is unseen and also compare unimodal versus multimodal training.
-5.1 Classification
-For classification, we utilize a classifier architecture comprising one or multiple pretrained encoder(s) , projection(s) , and a classification head consisting of 2 layers.
-We contrast three scenarios: In the baseline setting, all weights are initialized randomly, meaning the model lacks any pretrained knowledge.
-In the subsequent scenarios, the encoder and projection head is initialized with weights from our pretrained model.
-In one case, the weights of the encoder and projection head are fixed (frozen), whereas in the other case, we fine-tune them during the classifier’s training process.
-We train the model with all possible input modalities, namely accelerometer, poses, and videos, separately.
-Furthermore, we experiment with multimodal input by providing all three modalities to the model and concatenating the outputs from the respective projections before feeding them into .
-In addition, we train with different proportions of the training data, ranging from one to 100 percent, with the training data being sampled per class to ensure balanced representation.
-To guarantee comparability of all results, we train each classifier with the same settings: We utilize a learning rate of 0.0001 and training is conducted for a maximum of 200 epochs, incorporating early stopping after 25 epochs without improvement.
-We employ a weighted cross-entropy classification loss to address class imbalance, given by:
-where is the number of classes, is a one-hot encoded ground truth activity label and is the predicted label vector of for input , that is, the application of our classification layers to obtain .
-The weight for class is calculated by dividing the number of data samples of the largest class by the number of instances in class itself.
-To ensure validity of the results, we train 20 models independently for each configuration, with the training data being randomly selected anew for each run (for proportions less than 100%) and the non-pretrained weights being reinitialized randomly.
-5.1.1 FLAG3D
-The FLAG3D Tang et al.
-(2023) dataset is a large-scale 3D fitness activity collection with language instruction of 60 categories.
-The dataset comprises (a) 7,200 3D activity sequences performed by 10 individuals acquired from an advanced MoCap system, (b) 172,800 2D RGB videos rendered from the 3D MoCap data including different scenes, avatars, and camera positions, and (c) 7,200 authentic videos of ten persons captured by cost-effective smartphones in both indoor and outdoor natural settings.
-Additionally, FLAG3D offers comprehensive and meticulously crafted sentence-level language instructions for each fitness activity.
-We evaluate our method on all 7,200 real-world videos from Flag3D.
-To prepare the data, we follow the same approach as described in Section 3.2: First, we convert the videos to 50 fps.
-Then, we extract CLIP video features, ADA text embeddings, virtual sensor data, and poses (VideoPose3D) with IMUTube.
-We discard 943 videos for which IMUTube was unable to generate any sensor data.
-For the remaining 6,257 videos, we retrieve cumulatively over 21 hours of useful data.
-We split the data based on the person IDs performing the exercises into train (5 persons), validation (2 persons), and test (3 persons) sets.
-We apply a sliding window with a window size of 100 frames (2 seconds) and a step size of 50 frames (1 second).
-We also evaluate our method on a subset of 7,200 rendered videos, from which IMUTube was capable to extract sensor data for 6,832 videos (totaling 24 hours), processing them analogously to the real-world videos.
-5.1.2 MM-Fit
-The MM-Fit Strömbäck et al.
-(2020) dataset includes multi-view RGB-D video recordings, accompanied by 3D pose estimations extracted from single-view RGB frames with OpenPose and VideoPose3D.
-Additionally, it comprises a comprehensive compilation of time-synchronized inertial sensor data sourced from two smartphones, two smartwatches, and a earbud worn by participants during workout sessions.
-10 Participants performed 21 full-body workout routines in total while positioned in front of two depth cameras, with five smart devices strategically placed to collect inertial sensor data.
-The task is delimited to the recognition of exercises performed by individual participants.
-Thus, each workout session involves a solitary participant.
-This setup is tailored for home workout environments.
-Each workout session encompasses three sets of 10 exercises, with 10 repetitions for each exercise.
-The selected exercises encompass well-known resistance training routines, including squats, lunges (with dumbbells), alternating-arm bicep curls, sit-ups, push-ups, seated overhead dumbbell triceps extensions, standing dumbbell rows, jumping jacks, seated dumbbell shoulder press, and dumbbell lateral shoulder raises.
-We conduct the same preprocessing steps as for our dataset and FLAG3D but solely extract ADA text embeddings and CLIP video features.
-For pose and sensor, we utilize the 3D pose and Smartwatch accelerometer data of the left and right wrist provided by MM-Fit.
-Furthermore, we consider two scenarios: training the classifiers with and without including the “No Activity” category.
-Additionally, we calibrate sensor and pose data using a probability integral transform function Conover and Iman (1981), following the implementation of Kwon et al.
-(2021).
-In the unsupervised setting, all real data is calibrated with the virtual train data; whereas in the supervised case, label information is utilized to perform the calibration separately for each class.
-5.2 Evaluation of Generalization Performance on Unseen Datasets
-To evaluate performance in a more realistic scenario, we measure the model’s capability to match input modalities with the correct labels of the MM-Fit dataset in the scenario where the dataset is unseen.
-For that, we first calculate the ADA text embeddings for all labels of the MM-Fit dataset.
-For the “No Activity” class, which doesn’t correspond to a specific activity, we generate its embedding by taking the mean of the embeddings of all labels.
-For a given data sample of input modality , we then choose labels based on the closest similarity score by calculating for all labels .
-6 Results
-6.1 FLAG3D
-Table 4 demonstrates that the pretrained model consistently outperforms the baseline across all input modalities with respect to Macro -Score on both the real-world and the rendered dataset when utilizing either 2% or 100% of the training data.
-When comparing the baseline to our unfrozen pretrained model across both datasets and all types of input, we observe a absolute performance increase ranging from 0.042 (rendered dataset, video input) to 0.219 (real-world, multimodal) with just 2% of the training data.
-For the full dataset (100% of the training data), the improvement spans from 0.029 (real-world, video) to 0.208 (real-world, multimodal).
-This indicates that our pretrained encoders have learned a robust representation of the respective input modality, thereby enhancing classification performance even on a new dataset.
-We expect that our pretrained, unfrozen encoders would outperform the frozen ones by adapting more effectively to the training data.
-This expectation was confirmed in 7 out of 8 scenarios when using the full dataset for training, with a absolute performance increase of 0.014 on average.
-However, with only 2% of the training data, the frozen encoders show even superior performance in 3 out of 8 scenarios, and the unfrozen encoders do not perform significantly better on average, suggesting that the pretrained representation could not be enhanced with such a limited amount of training data.
-Generally, poses perform best out of all input modalities.
-While the multimodal setting does not show substantial improvement for the rendered dataset, a slight enhancement (0.015 for 2% train and 0.028 for 100% train) is observed for our pretrained, unfrozen model in the real-world setting when concatenating all modalities, compared to just using poses as input.
-From the boxplots presented in Figure 3, which compare the performance of the models across all modalities and various training fractions ranging from 1% to 100%, we observe that the pretrained models significantly outperform the baseline for all training fractions and most modalities.
-The only exception is the video modality, which performed worse out of all modalities and only slightly better than the baseline.
-However, with a limited amount of training data, our model demonstrates better performance relative to the baseline in the classification task over all modalities.
-6.2 MM-Fit
-Table 4 presents the results of experiments conducted with and without the inclusion of the “No Activity” class, across all input modalities and calibration configurations (none, unsupervised, and supervised), for both 2% and 100% training data fractions.
-We find that applying supervised, per class calibration generally enhances performance, whereas unsupervised calibration tends to diminish performance for sensor input.
-For pose data and multimodal inputs (which include poses), our pretrained encoders are unable to yield results without calibration—the training loss failed to decrease, remaining at infinity for 25 epochs.
-Notably, both IMUTube and MM-Fit utilize VidePose3D for pose extraction, but IMUTube incorporates extensive additional pre- and post-processing steps.
-Given that our pretraining utilized these poses generated by IMUTube, we speculate that our pose representation learned is significantly different from that of the MM-Fit poses.
-This difference likely hinders the model’s ability to learn when no calibration is performed.
-Indeed, calibration rectifies this issue.
-Excluding the scenarios without calibration for pose and multimodal inputs, our method surpasses the baseline in almost every case.
-The only exception is observed when training with sensor input and 100% training data, excluding the “No Activity” class and applying supervised calibration, where the baseline achieves a slightly better result (0.822 versus 0.798), which is due to the fact that MM-Fit uses real, rather than simulated sensor data.
-Leveraging multimodal inputs (i.e., concatenating all modalities) achieves the highest performance, suggesting that HAR benefits from the diverse perspectives of multiple modalities.
-In this multimodal scenario, our method, compared to the baseline, improves the Macro -Score from 0.945 to 0.965 (with “No Activity”) and from 0.973 to 0.999 (without “No Activity”) when trained on the full dataset.
-Furthermore, with only 2% of the training data available, we improve performance from 0.744 to 0.932 (with “No Activity”) and from 0.819 to 0.992 (without “No Activity”).
-These findings show that our model has significantly better performance compared to the baseline and learns more efficiently when only a small amount of training data is available.
-Figure 4 illustrates boxplots that exemplify the classification performances on accelerometer data input for all training fractions.
-It shows that our model significantly outperforms the baseline when trained with a minimal amount of training data, and generally, we observe a notable reduction in variance for our pretrained models compared to the baseline.
-This suggests that our pretrained sensor encoder has effectively learned a meaningful representation of real accelerometer data.
-Table 6.2 demonstrates the generalization performance of our representation on MM-Fit in the scenario where the dataset is unseen.
-In this analysis, we contrast the outcomes of unimodal training—where models were individually trained to align sensor-text, pose-text, and video-text—with those of multimodal training, where we tasked a single model to align all four modalities concurrently.
-Our findings reveal that, after implementing supervised calibration for pose and sensor modalities, our model surpasses random chance performance in all but one scenario.
-The peak top-1 performance achieves a Macro -Score of 0.638 for pose input.
-The only exception is the accelerometer input when the “No Activity” category is included.
-In this case, the top-1 performance is for unimodal and for multimodal training, both of wich are less than .
-This is also the only instance where multimodal training does not yield a performance enhancement, being outperformed by a small margin of 0.005.
-Conversely, in all other training configurations, multimodal training demonstrates superior performance over unimodal training.
-On average, multimodal training increases the performance by 0.069 (top-1).
-This supports our narrative that multimodal training can augment the model’s comprehension of various inputs.
-|c|c|c|ccc|ccc|
-Input Modality Calibration Macro -Score
-Unimodal Training Multimodal Training
-Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
-With “No Act.”
-Pose No
-Unsupervised
-Supervised
-Video No
-Sensor No
-Unsupervised
-Supervised
-Without “No Act.”
-Pose No
-Unsupervised
-Supervised
-Video No
-Sensor No
-Unsupervised
-Supervised
-7 Conclusion
-In this study, we have developed a novel approach to HAR by introducing a joint feature space that integrates video, pose, simulated accelerometer, and textual data.
-Our model was trained on a large, self-created multimodal fitness dataset containing over 78 hours of recordings.
-We evaluated our method using the FLAG3D dataset and the MM-Fit dataset.
-On MM-Fit, our model achieved an impressive Macro -Score of up to 0.992 with only 2% of the train data and 0.999 when all available training data was used.
-Moreover, in the scenario where the MM-Fit dataset was unseen, we demonstrated a generalization performance of up to 0.638.
-Our results reveal that our pretrained encoders effectively capture a meaningful representation for recognizing human activities.
-Furthermore, our approach not only demonstrates impressive results on an unseen dataset but also is capable of enhancing performance when applied to real-world sensor data.
+6.3 Zero-Shot
+Table 6.3 demonstrates zero-shot results of MuJo on all four datasets.
+For MM-Fit and MotionSense, multimodal training outperforms unimodal training, achieving constantly higher top-1, top-3, and top-5 accuracy and Macro -Scores.
+The MHEALTH dataset shows minimal differences between both training methods.
+However, for the MYOGYM dataset, which contains 30 gym exercises where all but 2 require equipment like machines, barbells, and more, mutimodal training performs worse.
+The reason for this might be that our pre-training was conducted on YouTube videos of fitness trainers primarily performing HIIT workouts, at home and not in the gym, focusing on bodyweight exercises.
+This domain shift may explain the decrease in performance from unimodal to multimodal training on MYOGYM: since the video features in our pre-training dataset lack representations of equipment-based exercises, they provide no valuable information for such types of exercises.
+Thus, the multimodal model might struggle more to recognize activities involving gym equipment.
+Consequently, unimodal training performs slightly better on this complex dataset, suggesting that target domain differences significantly impact performance.
+The overall low scores are mainly due to data divergences between our pre-training data and the evaluation datasets, particularly for pose (MM-Fit) and sensor data.
+Since we pre-trained with virtual sensor data and poses extracted by IMUTube, our models struggle when applied to real sensor datasets and variations in pose representation (MM-Fit), leading to decreased performance.
+Video features perform the best out of all modalities because they were extracted in the same way during pre-training.
+Further, the models were not pre-trained with the NULL class, which is represented as the mean of all target class embeddings.
+Introducing this unknown class during evaluation increases classification complexity and leads to an additional drop in performance.
+|c|c|c|ccc|ccc|ccc|ccc|
+Dataset
+With
+“NULL”
+Input Modality Metrics
+Accuracy Macro -Score
+Unimodal Training Multimodal Training Unimodal Training Multimodal Training
+Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
+MM-Fit ✗ Sensor .112 .310 .491 .169 .378 .575 .084 .274 .461 .114 .341 .536
+Pose .159 .369 .587 .196 .561 .671 .027 .233 .443 .061 .420 .574
+Video .393 .565 .652 .501 .660 .733 .285 .494 .611 .427 .616 .712
+✓ Sensor .027 .396 .862 .064 .559 .879 .027 .165 .400 .036 .227 .498
+Pose .035 .080 .127 .044 .282 .409 .006 .188 .370 .015 .301 .475
+Video .073 .113 .454 .111 .477 .920 .107 .240 .395 .201 .417 .680
+MHEALTH ✗ Sensor .117 .317 .515 .100 .335 .536 .093 .268 .454 .081 .273 .484
+✓ Sensor .161 .584 .817 .164 .638 .829 .054 .213 .418 .052 .203 .417
+MYOGYM ✗ Sensor .083 .161 .235 .052 .122 .172 .054 .117 .190 .032 .093 .135
+✓ Sensor .021 .046 .147 .009 .026 .070 .021 .054 .113 .005 .041 .078
+MOTIONSENSE ✗ Sensor .143 .428 .765 .146 .481 .867 .076 .353 .745 .104 .453 .850
+7 Conclusion and Future Work
+In this study, we introduced FiMAD, a multimodal dataset that contains over 78 hours of recordings.
+Our method, MuJo, is a novel pre-training approach that learns a joint feature space integrating video, pose, simulated accelerometer, and textual data.
+Our approach provides gains in real datasets such as MM-Fit, MyoGym, MotionSense, and MHEALTH, showing that our pre-training method captures a meaningful representation for HAR.
+Despite being trained solely on virtual IMU data, it enhances performance on real accelerometer data, especially when less data was used for training.
+Evenmore, our approach offers the flexibility to be applied not only to sensor inputs but also to video and pose inputs.
+Finally, although our zero-shot results are not outstanding, they demonstrate that multimodal training can enhance performance compared to unimodal training, if the target domain does not differ too much from our pre-training domain.
+While our study presents a robust and innovative approach to HAR by integrating multiple data modalities into a joint feature space, there are several avenues for future research and enhancement.
+One potential improvement is to fine-tune the text and vision models during pre-training, rather than using pre-calculated text and video embeddings.
+Additionally, we plan to conduct a more detailed evaluation in the future, including fine-tuning MuJo on the target dataset using real sensor data, rather than relying solely on our pre-trained model.
+Furthermore, we aim to enhance the handling of the NULL class, which can be achieved by using a large language model to label the chapters of FiMAD that do not contain relevant activities as NULL.
+This would provide a clear representation for the NULL class in our dataset.
+Future work also includes incorporating other proxy tasks in our framework during pre-training as well as scaling the networks.
+In that setup we aim to train a single MuJo model instead of training a separate one for each target dataset when the sensor positions differ.